Prepare 2 L of a 35% alcohol solution using a 70% alcohol and a 15% alcohol.
Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

Prepare 2 L of a 35% alcohol solution using a 70% alcohol and a 15% alcohol.
-----------
Do it like this:
if a 14% acid solution is mixed with a 19% acid solution to produce 45 gallons of a 15% acid solution how much (in gallons) of each solution should be used?
--------------
f = 14%
n = 19%
----
f + n = 45 (total gallons)
14f + 19n = 15*45 (total acid)
-----
n = 45 - f
Sub for n
14f + 19*(45-f) = 675
14f + 855 - 19f = 675
-5f = -180
f = 36 gallons of 14%
n = 9 gallons of 19%

Answer by ewatrrr(24785)

 
Hi,
Prepare 2 L of a 35% alcohol solution using a 70% alcohol and a 15% alcohol
alcohol = alcohol
.70x + .15(2L-x) = .35*2L
.55x = .70-.30
x = .40/.55 = .73L of 70%alcohol and 1.27L of 15%alcohol